the perimeter of a rectangle is 102 inches , and the length of the diagonal is 39 inches. find the dimensions of the rectangle
Answer by tommyt3rd(5050)

1)
P=2(l+w)=102
l+w=51

2) l^2+w^2=39^2

using substitution...
(l-51)^2+l^2=1521
l={15, 36}
so the dimensions are 15in x 36in